# RocketKV: Accelerating Long-Context LLM Inference via Two-Stage KV Cache Compression

### v1

@@ 3.4 Second Stage: Hybrid Attention
Previous methods on dynamic KV token selection often leverage the sparsity in the head or sequence dimension of the Q or K tensor [31, 25, 28] to estimate top-k attention scores. Unfortunately, relying on the sparsity of a single dimension can only achieve a certain degree of compression ratio, beyond which the accuracy could drop rapidly, as shown previously in Figure 1. In contrast, we propose a hybrid attention method which takes advantage of sparsity in both head and sequence dimensions together to achieve better estimation accuracy on KV token indices with top-k attention scores.
@@
Figure 3 shows the detailed implementation of our proposed algorithm, which is inspired by Quest [31] and SparQ [25]. Our hybrid attention algorithm can be decomposed into three steps:
@@
  • Step 1: Group tokens in key tensor into consecutive pages along the sequence dimension and store element-wise min and max of each page as auxiliary storage. They are stored along the head dimension to enable efficient gathering in Step 2. The auxiliary storage is updated accordingly each time a new key token is generated and stored.
@@
  • Step 2: For each query q, accumulate $|q|$ in group dimension and find the r largest indices along the head dimension. Then fetch only the corresponding indices in either element-wise min or max tensors depending on the sign of the sum of $q$ in group dimension at those indices. Finally calculate approximation attention scores based on those partial $q$ and element-wise vectors, and find top-k indices along the sequence dimension.
@@
  • Step 3: Perform sparse attention by fetching the full key and value vectors from the top-k indices.
In our hybrid attention algorithm, the $K$ tensor is compressed along the sequence dimension first for auxiliary storage, and then selectively fetched along the head dimension for approximation attention calculation. For a given compression ratio $c$, we evenly split it between those two steps so that each step results in a compression ratio of $\sqrt{c}$. This would avoid significant accuracy loss caused by over-compression on a single dimension. Our method is fully compatible with GQA because all selections on the key and value vectors are based on per attention group rather than attention head, similar to our enhancement in SnapKV++.
@@
Table 1: Normalized KV cache storage (including auxiliary data) and traffic comparison between RocketKV and other methods.
@@
RocketKV $c$ $1/\sqrt{c}+2/c^{3/4}$ $1/c$
3.5 Cost Analysis and Comparison
The decode phase of LLM inference is typically memory bound [25] so the time spent in the attention module is roughly proportional to the total memory traffic. In this work, we use token budget $t$ to estimate the amount of memory traffic for each attention operation in the decode phase (we mainly focus on KV cache traffic since it contributes to the majority of memory traffic in this scenario). For example, a token budget of 512 means each attention module needs to fetch an equivalent total amount of 512 key and value pairs from the memory. For models with GQA, this token budget is defined for the entire attention group rather than each attention head. For a given sequence length $S$, the total compression ratio $c$ can be defined as $c=S/t$.
In RocketKV, the total compression ratio is evenly split between two stages with each stage having a compression ratio of $\sqrt{c}$. Therefore, in the first stage, SnapKV++ results in both KV cache storage and traffic reduction of $\sqrt{c}$. In the second stage, we need to take the additional memory storage overhead introduced by approximation attention into consideration. Since we further evenly split the compression ratio $\sqrt{c}$ into $c^{1/4}$ between step 1 and 2 in the hybrid attention algorithm, it introduces a memory storage overhead of $(1/\sqrt{c})\times(1/c^{1/4})\times 2=2/c^{3/4}$ where $1/\sqrt{c}$ is the relative KV cache storage after SnapKV++ and both element-wise min and max tensors will introduce a storage overhead of $1/c^{1/4}$ on top of it. Therefore, the total KV cache storage and traffic in RocketKV are $1/\sqrt{c}+2/c^{3/4}$ and $1/c$ of the full KV baseline, respectively. Table 1 compares the KV cache storage and traffic of RocketKV against other methods at a given compression ratio $c$ and we can see that while all methods lead to KV cache traffic savings, only RocketKV and SnapKV provide additional KV cache storage savings but Quest and SparQ require extra storage for auxiliary data.
In terms of system-level optimizations, RocketKV is fully compatible with FlashAttention [7] as it does not modify attention in the prefill phase, as well as Tensor Parallelism [27] as all operations are symmetric across attention heads/groups.
@@ 4.1 Experimental Settings
Figure 4: Comparing the accuracy of RocketKV with other methods under LongBench ((a) to (c)), Needle-in-a-Haystack ((d) to (f)), and RULER with various sequence lengths ((g) to (r)).
We conduct our experiments on three widely used long-context models: Llama3.1-8B-Instruct [1], Mistral-7B-Instruct-v0.2 [13], and LongChat-7B-v1.5-32k [19]. We use the notations Llama3.1-8B-Ins, Mistral-7B-Ins-v0.2, LongChat-7B-v1.5 for them, respectively throughout the paper. The first two models are based on GQA, while the last one is based on MHA. For downstream tasks, we utilize LongBench [4], Needle-in-a-Haystack [16] and RULER [11] benchmarks. We compare RocketKV with several other methods: Full-KV, Exact-TopK, DuoAttention [35], SnapKV [20], Quest [31], and SparQ [25]. Exact-TopK serves as an oracle method for sparse attention with exact top-k KV token selection. We evaluate all methods across various KV token budgets per attention group except for Full-KV. RocketKV, Quest, and SparQ involve additional memory traffic for top-k approximation, which is converted into an equivalent KV token budget such that the total token budget precisely reflects the total amount of memory loads in the attention module. More details on our experimental settings can be found in Appendix A.
@@ 4.2 Accuracy Results
In our accuracy evaluation, we vary the token budget of each method from 256 to 4096 except for the Full-KV and then compare the average accuracy across all individual tasks for each benchmark. More detailed results with accuracy breakdown for individual tasks can be found in Appendix B.
@@
LongBench Benchmark: The first row in Figure 4 shows the average score comparison of RocketKV with other methods on LongBench across all three models. Based on the figure, we can see that RocketKV consistently outperforms all other methods, especially with the lower token budgets. For Llama3.1-8B-Ins, RocketKV achieves almost no accuracy loss with a token budget of 512 and above, and only 1.1% average accuracy drop with a token budget of 256. RocketKV results in slightly higher accuracy losses for Mistral-7B-Ins-v0.2 and LongChat-7B-v1.5, which might be because Llama3.1-8B-Ins is better-trained than the other two models, making it more robust to sparse attention methods, as a similar trend can be found with Exact-TopK. SparQ performs well on Llama3.1-8B-Ins and LongChat-7B-v1.5 with a token budget of 1024 and above but underperforms all other methods on Mistral-7B-Ins-v0.2. SnapKV achieves relatively good accuracy with a token budget of 1024 and above across all three models, however, it cannot further reduce the token budget to 512 or below for tasks in LongBench with 512 maximum generated tokens. This is because it only compresses the KV cache of the input prompt but not the generated tokens, so the token budget cannot reach 512 or below for those tasks. Although we cannot plot the average accuracy score of SnapKV for those cases, the accuracy scores for the remaining tasks are available in Appendix B.3, and we can still see widening accuracy gaps between SnapKV and RocketKV for those tasks.
@@
Needle-in-a-Haystack Benchmark: Presented in the second row of Figure 4, RocketKV achieves near the Full-KV accuracy across all models, even with a token budget of 512, and experiences only negligible accuracy loss with a token budget of 256. In contrast, all other methods suffer substantial accuracy drops. For example, the accuracy of SnapKV for Mistral-7B-Ins-v0.2 and LongChat-7B-v1.5 decreases by more than 20% compared to the Full-KV, even with a token budget of 1024. We present the heatmaps of RocketKV under various token budgets in Appendix B.2.
@@
RULER Benchmark: For the RULER benchmark, we evaluate all methods across three models with varying sequence lengths as shown in Figure 4 (rows 3-5). Again, RocketKV shows robust performance and a clear advantage over all other methods across various token budget and sequence length settings. Overall, we can see that the accuracy loss of RocketKV is negligible under short sequence lengths, and gradually becomes larger as sequence length increases. We believe this is because those models are less robust to sparse attention under longer sequence lengths. Notice that the accuracy gaps between RocketKV and other methods become even wider under longer sequence lengths.
(a) End-to-end speedup
@@
Figure 5: End-to-end speedup and peak memory savings of RocketKV with various token budgets compared to the Full-KV.
@@ 4.3 Efficiency Results
Our efficiency experiments are conducted with Llama3.1-8B-Ins [1] under FP16 precision, running on an NVIDIA H100 GPU at a batch size of 1. Similar to SparQ, we leverage gpt-fast [17], a low-latency Python-native
@@
LLM framework for running the efficiency experiments. We found a python-based implementation of RocketKV under gpt-fast is sufficient to demonstrate its efficiency benefit, but it could be further improved with customized CUDA kernels and more advanced frameworks such as FlashInfer [36]. Figure 5 demonstrates the end-to-end speedups and peak memory savings of RocketKV with varying token budgets at the decode phase, with all values normalized to the Full-KV. Notice that the peak memory usage measured in the figure includes memory allocations for weights, activations, KV cache, and all other metadata at the decode phase. As shown in the figure, RocketKV with token budgets of 512 and 256 achieves up to 2.9$\times$ and 3$\times$ end-to-end speedups as well as 30% and 31.4% peak memory savings, respectively. Additionally, increasing the sequence lengths results in almost linear improvements in both speedup and peak memory savings.
(a) LongBench, Llama3.1-8B-Ins
(b) LongBench, Mistral-7B-Ins-v0.2
(c) NIAH, Llama3.1-8B-Ins
(d) NIAH, Mistral-7B-Ins-v0.2
Figure 6: LongBench and Needle-in-a-Haystack accuracy comparison between SnapKV and SnapKV with GQA enhancement.
4.4 Ablation Study
In this subsection, we present a series of ablation studies to further demonstrate the effectiveness of RocketKV. More ablation results can be found in Appendix B.
4.4.1 SnapKV with GQA enhancement
As illustrated in Figures 6, SnapKV with the grouped-query attention (GQA) [2] enhancement consistently outperforms the original SnapKV, especially with low token budgets. This is because with the GQA enhancement, the compressed KV tokens are shared across the entire attention group which is far more effective than storing them separately within each attention head. Recall that, the GQA enhancement does not make any difference for LongChat-7B-v1.5 since it uses MHA in the attention module.
(a) Llama3.1-8B-Ins, SeqLen=16K
(b) Llama3.1-8B-Ins,SeqLen=32K
(c) Llama3.1-8B-Ins, SeqLen=64K
(d) Llama3.1-8B-Ins, SeqLen=96K
(e) Mistral-7B-Ins-v0.2, SeqLen=8K
(f) Mistral-7B-Ins-v0.2,SeqLen=16K
(g) Mistral-7B-Ins-v0.2,SeqLen=24K
(h) Mistral-7B-Ins-v0.2,SeqLen=32K
(i) LongChat-7B-v1.5, SeqLen=8K
(j) LongChat-7B-v1.5, SeqLen=16K
(k) LongChat-7B-v1.5, SeqLen=24K
(l) LongChat-7B-v1.5, SeqLen=32K
Figure 7: Comparing RocketKV with different kernel sizes (7, 63, and 511) under RULER with various sequence lengths.
4.4.2 RocketKV kernel size selection
We run extensive experiments to examine the impact of the kernel sizes for pooling on the overall accuracy of RocketKV. We primarily focus on the RULER benchmark to understand the optimal kernel size selection across different sequence lengths, as shown in Figures 7. For Llama3.1-8B-Ins, a kernel size of 63 performs best for sequence lengths of 16K and 32K, while a kernel size of 511 performs best for sequence lengths of 64K and 96K. Therefore, we simply set the switching threshold to 48K as the middle point between 32K and 64K. For Mistral-7B-Ins-v0.2, a kernel size of 63 generally performs best across all sequence lengths so there is no need to switch to other kernel sizes. For LongChat-7B-v1.5, a kernel size of 63 performs best for a sequence length of 8K, and a kernel size of 511 performs best for remaining sequence lengths; so, we set the switching threshold to 12K. We found adaptive kernel selection with empirically determined kernel sizes and threshold generalizes well in other benchmarks (shown in Appendix B.1). Hence, we do not explore more sophisticated solutions.
(a) LongBench, Llama3.1-8B-Ins
(b) LongBench, Mistral-7B-Ins-v0.2
(c) LongBench, LongChat-7B-v1.5
(d) NIAH, Llama3.1-8B-Ins
(e) NIAH, Mistral-7B-Ins-v0.2
(f) NIAH, LongChat-7B-v1.5
Figure 8: LongBench and Needle-in-a-Haystack accuracy comparison among Hybrid, Quest, and SparQ.
4.4.3 Comparing Hybrid, Quest, and SparQ
To illustrate the effectiveness of hybrid attention, we compare the accuracy of the standalone hybrid attention mechanism (Hybrid) against Quest and SparQ in the Figures 8. In all cases, hybrid attention consistently outperforms Quest and SparQ, especially at low token budgets. This clearly demonstrates the advantage of hybrid attention which intelligently leverages approximations in both head and sequence dimensions compared to single dimension approximation methods such as Quest and SparQ.

### v2

@@ 3.4 Second Stage: Hybrid Sparse Attention
Previous methods on dynamic KV token selection often estimate top-k KV indices with reduced computation along a single dimension. For example, Quest (Tang et al., 2024) uses element-wise minimum and maximum values to represent continuous pages along the sequence dimension. Meanwhile, SparQ (Ribar et al., 2024) and Loki (Singhania et al., 2024) leverage sparsity in the head dimension to conduct low-rank estimations. Unfortunately, relying on one-dimensional sparsity can only achieve a certain degree of compression ratio, beyond which the accuracy could drop rapidly, as shown previously in Figure 1. In contrast, we propose hybrid sparse attention (HSA), which takes advantage of two-dimensional reduction, in both sequence and head dimensions together to achieve better estimation accuracy on KV token indices with top-k attention scores.
@@
Figure 4 shows the detailed implementation of our proposed algorithm, which is inspired by Quest (Tang et al., 2024) and SparQ (Ribar et al., 2024). Our HSA algorithm can be decomposed into three steps:
@@
  • Step 1: Group tokens in key tensor into consecutive pages along the sequence dimension and store element-wise maximum $(K_{\max})$ and minimum $(K_{\min})$ values of each page as auxiliary storage similar to Quest. Unlike Quest, they are stored with a different layout by aligning along the head dimension to enable efficient gathering in Step 2. The auxiliary storage is updated accordingly each time a new key token is generated.
@@
  • Step 2: For each query $q$, find $k_{1}$ largest absolute values along the head dimension. Then, fetch only the corresponding indices in either element-wise maximum or minimum tensors, depending on the sign of $q$ at those indices. The goal is to compute element-wise $\max(q\times K_{\max},\ q\times K_{\min})$ for each page to approximate the highest possible attention scores within a page. To further reduce the approximation overhead, we only calculate on $k_{1}$ partial positions along the head dimension with a large magnitude of $q$ and ignore others, similar to SparQ. Once the approximation attention scores are calculated, $k_{2}$ indices with the largest attention scores along the sequence dimension are selected.
@@
  • Step 3: Perform sparse attention by fetching the original key and value vectors from the predicted $k_{2}$ indices.
Our HSA algorithm is fully compatible with GQA. To achieve this, we perform all key tensor selections on a per attention group basis. More details are shown in Algorithm 1 where we perform the sum of $q$ or $|q|$ in the group dimension as needed to guarantee that all attention heads within a group are making the same selection at each step.
3.5 RocketKV-MT
In a multi-turn conversation setting, KV tokens that were pruned in earlier turns may become essential for answering queries in later turns because KV token importance could vary significantly between different turns  (Li et al., 2025). Consequently, permanently removing these tokens could lead to a noticeable accuracy drop in subsequent turns.
To mitigate this issue, we introduce a multi-turn variant of our approach called RocketKV-MT. The core idea in RocketKV-MT is to avoid permanently evicting any KV tokens during the first stage; instead, all KV tokens are retained in memory across turns, ensuring that no potentially useful context is lost. Meanwhile, to preserve computational efficiency, the second stage still performs dynamic selection over the subset of KV tokens filtered by the first stage, similar to the original RocketKV. In other words, the model generates responses using a reduced KV token set for speed, while reserving the full KV history for future turns. For example, suppose the first stage of RocketKV-MT retains only $N$ out of $M$ total KV tokens from the input prompt in the first turn. RocketKV-MT will still keep all $M$ tokens in memory but restrict the second stage to dynamically select from these $N$ input tokens (plus any new generated tokens) during the decode phase. In the next turn, the full set of previously stored KV tokens (all $M$ input tokens plus all output tokens) is added to the new input KV cache. The filtering process (with SnapKV in our case) is applied again on this whole input KV cache to select a new subset of important tokens for this turn’s decode phase. By following this strategy, RocketKV-MT achieves similar decoding speedups to RocketKV in each turn, while retaining the full KV cache history across all turns. This approach effectively eliminates the accuracy degradation caused by permanent KV token eviction in multi-turn scenarios at the cost of no memory storage savings.
3.6 Adaptive Compression Decomposition
The decode phase of LLM inference is typically memory bound (Ribar et al., 2024); so, the time spent in the attention module is roughly proportional to the total memory traffic. In this work, we use the token budget $t$ to estimate the amount of memory traffic for each attention operation in the decode phase (we mainly focus on KV cache traffic since it contributes to the majority of memory traffic in this scenario). For example, a token budget of 512 means each attention module needs to fetch an equivalent total amount of 512 key and value pairs from memory. Unlike prior work (Tang et al., 2024) where the token budget only reflects the memory traffic for top-k attention (Step 3 in HSA), we define the token budget to also include the memory traffic of  top-k estimation (Step 2 in HSA). By doing this, the token budget can reflect the overall memory traffic in the attention module more precisely. For simplicity, we evenly split the token budget between these two steps for HSA and all other dynamic KV token selection methods in our later experiments. For models with GQA, this token budget is defined for the entire attention group rather than each attention head. For a given sequence length of $S$, the total compression ratio of $c$ can be defined as $c=S/t$.
Since RocketKV is a two-stage KV cache compression framework where the filtered KV token set by the first stage serves as the input of the second stage for dynamic selection, it is important to determine the intermediate token budget for the filtered KV token set. For an overall compression ratio of $c$, we define a split factor $r$ so that $c$ is split into $c^{r}$ for the first stage and $c^{(1-r)}$ for the second stage, where $10<=r<=$ 1. We use the following formula to adaptively determine $r$:
$r=min(0.2+0.06*log_{2}(c),0.8)$
The insight behind this formula is that when $c$ is small, we would like to minimize the number of permanent evicted KV tokens in the first stage to prevent information loss. As $c$ increases, the accuracy drop caused by HSA in the second stage gets larger thus it is better to assign a higher compression ratio to the first stage because SnapKV can estimate important KV tokens more precisely using exact attention scores. We further limit the range of $r$ between 0.2 and 0.8 to balance the compression decomposition between these two stages.
For compression decomposition within HSA, we simply split the compression ratio $c^{(1-r)}$ evenly between the sequence and head dimensions so that each dimension gets a compression ratio of $c^{(1-r)/2}$. Notice that the compression ratio along the sequence dimension is equivalent to the page size, so we need to round it up the nearest integer $\lceil c^{(1-r)/2}\rceil$ and the other dimension gets $c^{(1-r)}/\lceil c^{(1-r)/2}\rceil$.
Because RocketKV strategically decomposes KV cache compression across multiple stages and dimensions, it significantly enhances the potential for high compression ratios while maintaining strong accuracy—surpassing methods that rely on single-stage, single-dimension approaches. For example, given a compression ratio of 64×64$\times$64 ×, the split factor can be calculated as $r=0.2+0.06*log_{2}(64)=0.56$. Therefore, the compression ratio is split into $64^{0.56}=10.3\times$ in the first stage and $64^{(1-0.56)}=6.2\times$ in the second stage. HSA further splits its compression ratio into 3×3$\times$3 × in the sequence dimension (with a page size of 3) and 2.1×2.1$\times$2.1 × in the head dimension. We can see that each individual stage and dimension gets assigned with a much smaller compression ratio after decomposition.
@@ 3.6 Adaptive Compression Decomposition
Table 1: Normalized KV cache storage (including auxiliary data) and traffic comparison between RocketKV, RocketKV-MT and other methods.
@@
RocketKV $c$ $1/c^{r}+2/c^{(1+r)/2}$ $1/c$
RocketKV-MT $c$ $1+2/c^{(1+r)/2}$ $1/c$
In RocketKV, the first stage results in both KV cache storage and traffic reduction of $c^{r}$. In the second stage, we need to take the additional memory storage overhead introduced by the approximation attention into consideration. Since we evenly split the compression ratio $c^{(1-r)}$ into $c^{(1-r)/2}$ (ignore the round up operation here for simplicity) between two dimensions in HSA, it introduces a memory storage overhead of $(1/c^{r})\times(1/c^{(1-r)/2})\times 2=2/c^{(1+r)/2}$ where $1/c^{r}$ is the relative KV cache storage after the first stage and both element-wise maximum and minimum tensors introduce a storage overhead of $1/c^{(1-r)/2}$ on top of it. Therefore, the total KV cache storage and traffic in RocketKV are $1/c^{r}+2/c^{(1+r)/2}$ and $1/c$ of the full KV baseline, respectively. RocketKV-MT does not lead to storage saving in the first stage, so its relative KV cache storage is $1+2/c^{(1+r)/2}$ instead. Table 1 compares the KV cache storage and traffic of RocketKV and RocketKV-MT against other methods at a given compression ratio $c$. We can see that while all methods lead to the same KV cache traffic savings, only RocketKV, DuoAttention and SnapKV provide additional KV cache storage savings, but RocketKV-MT, Quest and SparQ require extra storage for auxiliary data.
3.7 System Implications
RocketKV is fully compatible with FlashAttention (Dao et al., 2022) because it does not modify attention in the prefill phase. Additionally, it seamlessly integrates with tensor parallelism (Shoeybi et al., 2019) because all operations are symmetric across attention heads/groups. It is worth noting that both RocketKV and RocketKV-MT work well with the disaggregated serving system, where different GPUs are used for prefill and decode phases (NVIDIA, 2025; Qin et al., 2025). For RocketKV, the KV cache storage is reduced on both the prefill and decode GPU, as well as the KV cache transfer traffic between them. While the full KV cache needs to be stored in the prefill GPU for RocketKV-MT, only the filtered set needs to be transferred and stored in the decode GPU, resulting in the same communication and decoding benefits as RocketKV.
@@ 4.1 Experimental Settings
Figure 5: Comparing the accuracy of RocketKV with other methods on LongBench (a-c), Needle-in-a-Haystack (NIAH) (d-f), and RULER with various sequence lengths (g-r).
We conduct our experiments on three widely used long-context models: Llama3.1-8B-Instruct (AI@Meta, 2024), Mistral-7B-Instruct-v0.2 (Jiang et al., 2023), and LongChat-7B-v1.5-32k (Li et al., 2023). We refer to these models as Llama3.1-8B-Ins, Mistral-7B-Ins-v0.2, LongChat-7B-v1.5, respectively, throughout the paper. The first two models are based on GQA, while the last one is based on MHA. For downstream tasks, we utilize LongBench (Bai et al., 2024), Needle-in-a-Haystack (Kamradt, 2023), RULER (Hsieh et al., 2024), and SCBench (Li et al., 2025). We compare RocketKV and RocketKV-MT with several other methods: Full-KV, Exact-TopK, DuoAttention (Xiao et al., 2025), SnapKV (Li et al., 2024), Quest (Tang et al., 2024), and SparQ (Ribar et al., 2024). Exact-TopK serves as an oracle method for sparse attention with exact top-k KV token selection. Notice that RocketKV-MT acts the same as RocketKV in single-turn scenarios, so we only evaluate it on SCBench under multi-turn mode. We evaluate all methods across various KV token budgets per attention group, except for Full-KV. RocketKV, RocketKV-MT, Quest, and SparQ involve additional memory traffic for top-k approximation, which is converted into an equivalent KV token budget such that the total token budget precisely reflects the total amount of memory traffic in the attention module. More details on our experimental settings are available in Appendix A. We also conduct additional ablation studies in Appendix B.1.
@@ 4.2 Accuracy Results
In our accuracy evaluation, we vary the token budget of each method from 256 to 4096 for all single-turn benchmarks and from 1024 to 16384 for SCBench under multi-turn mode. We compare the average accuracy across all individual tasks for each benchmark. More detailed results with accuracy breakdown for individual tasks are available in Appendix B.
@@
LongBench Benchmark: The first row in Figure 5 shows the average score comparison of RocketKV with other methods on LongBench across all three models. Based on the figure, we can see that RocketKV consistently outperforms all other methods, especially under lower token budgets. For Llama3.1-8B-Ins, RocketKV achieves almost no accuracy loss with a token budget of 512 and above, and only 1.1% average accuracy drop with a token budget of 256. RocketKV results in slightly higher accuracy losses for Mistral-7B-Ins-v0.2 and LongChat-7B-v1.5, which might be because Llama3.1-8B-Ins is better-trained than the other two models, making it more robust to sparse attention methods, as a similar trend can be found with Exact-TopK. SparQ performs well on Llama3.1-8B-Ins and LongChat-7B-v1.5 with a token budget of 1024 and above but underperforms all other methods on Mistral-7B-Ins-v0.2. SnapKV alone achieves relatively good accuracy under low token budgets, but we still observe widening accuracy gaps between SnapKV and RocketKV as the token budget decreases.
@@
Needle-in-a-Haystack (NIAH) Benchmark: Presented in the second row of Figure 5, RocketKV achieves near the Full-KV accuracy across all models, even with a token budget of 256. In fact, it reaches 100% accuracy on Llama3.1-8B-Ins under 256 token budget, which corresponds to a compression ratio of over 400$\times$ with a maximum sequence length of 109K tokens. In contrast, all other methods suffer from substantial accuracy drops. While SnapKV leads to comparable accuracy to RocketKV on Llama3.1-8B-Ins, its accuracy for Mistral-7B-Ins-v0.2 and LongChat-7B-v1.5 decreases by more than 20% compared to Full-KV, even with a token budget of 1024. We present the heatmaps of RocketKV under various token budgets in Appendix B.2.
@@
RULER Benchmark: For the RULER benchmark, we evaluate all methods across three models with varying sequence lengths as shown in Figure 5 (rows 3-5). Again, RocketKV shows robust performance and a clear advantage over all other methods across various token budget and sequence length settings. Overall, we can see that the accuracy loss of RocketKV is negligible under short sequence lengths and gradually becomes larger as sequence length increases. We believe this is because these models are less robust to sparse attention beyond their effective sequence lengths as defined in the RULER paper (Hsieh et al., 2024). Notice that the accuracy gaps between RocketKV and other methods become even wider under longer sequence lengths.
Figure 6: Comparing the accuracy of RocketKV and RocketKV-MT with other methods on Llama3.1-8B-Ins for SCBench.
SCBench Benchmark: Figure 6 presents the accuracy comparison of various methods on SCBench under the multi-turn setting. While RocketKV still outperforms other methods under low token budgets, it provides lower accuracy than SparQ for token budgets of 8192 and beyond. Moreover, there is still a noticeable accuracy gap between RocketKV and Exact-TopK under all token budgets, showing room for further improvement. As we discussed earlier, this is primarily caused by the fact that KV token importance varies significantly across different turns, so the unimportant KV tokens evicted by earlier turns could lead to a significant accuracy drop in later turns. After fixing this issue in RocketKV-MT, we can see that RocketKV-MT achieves a significant accuracy boost with comparable accuracy to Exact-TopK across all token budgets.
(a) End-to-end speedup on A100
(a) End-to-end speedup on H100
@@
Figure 7: End-to-end speedup and peak memory savings of RocketKV with various token budgets compared to Full-KV.
@@ 4.3 Efficiency Results
Our efficiency experiments are conducted with Llama3.1-8B-Ins (AI@Meta, 2024) under FP16 precision, running on NVIDIA H100 and A100 GPUs at a batch size of 1. Similar to SparQ, we leverage gpt-fast (PyTorchLabs, 2023), a low-latency python-native LLM framework for running the efficiency experiments. We found a python-based implementation of RocketKV under gpt-fast is sufficient to demonstrate its efficiency benefit, but it could be further improved with customized CUDA kernels and more advanced frameworks such as FlashInfer (Ye et al., 2025). Figure 7 demonstrates the end-to-end speedups and peak memory savings of RocketKV with varying token budgets at the decode phase, with all values normalized to Full-KV. Notice that the peak memory usage is the same regardless of the underlying GPU, and it is measured includes memory allocations for weights, activations, KV cache, and all other metadata at the decode phase.
As shown in the Figure 7, RocketKV achieves up to 3.7$\times$ and 3.3$\times$ end-to-end speedups on an A100 and H100 GPU, respectively, as well as up to 32.6% peak memory saving. The maximum speedup on A100 is 12% higher than H100 because A100 has a lower memory bandwidth to compute ratio compared to H100. As a result, LLM inference execution is more memory-bound on A100 and can benefit more from the memory traffic savings of the KV cache offered by RocketKV. We expect the speedup of RocketKV will be even higher on cheaper GPUs such as RTX 4090/5090 since they are not equipped with High Bandwidth Memory (HBM).